November 8, 2013 Installation

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Friday, November 8, 2013 at 1:30 PM At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone (supported on request only) dial:
  • You can look up toll-free number at
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#

Desktop Sharing

You can view the computer desktop in the meeting room at JLab via the web.

  1. Go to
  2. In the "join a meeting" box enter the Hall D code: 3421244
  3. Fill in the participant registration form.

Up to GlueX/Hall-D Installation Meetings

Action Items From Last Meeting


  1. Announcements
  2. Updates to the Mechanical Installation Schedule ( Nov 8 Rendition of Install Schedule) (Tim)
  3. Electronics Installation Schedule (Fernando) [| Schedule]
    • Readiness of controls
  4. Beam Commissioning Plan Reviews Sequence (Elton)
  5. Installation Updates (Not a construction report)
    • Counting House Infrastructure
    • Trigger and DAQ Installation
    • Slow Control Installation
    • Tagger Magnet Installation
    • BCAL Installation
    • FCAL Installation
    • FDC Installation
    • CDC Installation
    • TOF Installation
    • Start Counter Installation
    • Microscope Installation
    • Tagger Array Installation
    • Pair Spectrometer Installation
  6. A.O.B.


Installation meeting:

CMU, Regina, FIU, FSU, JLab, IU:

Tim, schedule: Tagger magnet mapping is getting pushed back. Paul hopes for TOF installation earlier in December than pictured.

Fernando, electronics: AC power is on the downstream platform is in place. Grounding is now underway. Power should be available on Monday morning. Power for BCAL is in place but locked out. Cables are prepared for FDC -- crates and racks and hardware are in place for FDC testing. Network equipment will be moving to UPS power.

Elton: Presented the various steps and safety reviews needed to commission the beam going from Electron Beam to the tagger dump, to photon beam in the hall, to experiential equipment in Hall D.

Review of photon beam must happen at least 3 months prior to beam delivery. If it is scheduled for March, this means that inserting a radiator and putting photons in Hall D in April is not possible.

In addition each detector element needs a discussion of major operational and safety considerations. Emphasis should be put on preparing the TDR, as each system can provide a summary of the detector system that is sufficient for the review document.

Need to understand how the necessary reviews affect our commissioning schedule.

Installation Updates:

Hovanes (Counting House): work in counting house is slowed by RedHat licensing. Operating systems will be complete in a week or two.

Elliott (Networking): All purchased switches have been handed out for use in labs and need to reclaim for installation in the hall. Networking is ready for FCAL and will be ready for BCAL very soon. Most cable trays are installed. Fiber networking will take some time yet -- drawings are in preparation now.

Sasha (Trigger): Chris C. has measured links for fiber optic trigger cables. Trigger crates will be installed soon - maybe 2 weeks until power is available for trigger crates.

Elliott (Controls): Dave Butler continues to work on controls. Work is on schedule.

Tim (Tagger Magnet): Survey is happening today. There is a preliminary OSP. Hopefully mapping will happen at the end of next week.

Elton (BCAL): Cabling is complete. Manifolds installed and pressure tested. Electronics installation is underway.

John (FCAL): Plexiglas panels installed. LED boards need to be installed on the edges of plates. The FCAL darkroom control system works with the PLC. The next steps are to interface with the door magnets. The front cover has also been assembled. Hope to seal up darkroom in several weeks.

Lubomir (FDC): Survey indicates FDC packages are very flat and well centered. However, package is not rigid. Goal is to measure z position of inner FDC packages while they are inside the magnet. Survey reflectors have been made, but need to be installed. (Remove FDC, install reflectors, and reinstall FDC.) Chilled water will be tested in the floor before reinstallation in order to check for leaks. Hope that final installation will happen around end of November.

Naomi (CDC): CDC has been delivered to JLab. All wires have been checked and still OK. Two channels have been identified as problematic, but these are the same two that were identified as problematic at CMU.

Paul (TOF): Half of TOF has been delivered to JLab. 2 of 6 crates were opened: no indication of any movement or problem. Hope to start installation into planes in December.

Eric (Start Counter): Construction still underway.

Sasha (Microscope/Hodoscope): Aim for delivery at the end of the January, maybe several weeks late. Both will come to JLab at that time. Pair spectrometer counters will also be installed in January, on schedule.

Pair spectrometer magnet mapping: plan is for it take place in December after tagger magnet mapping is done.